Number of times the first card is face card and the second card is not face card


1
Expert's answer
2022-02-23T13:03:36-0500

Face cards is: jack, queen, king

not face cards is all others

Number of times when first card is face is 4 * 3 = 12

Number of times when second card is not face is 52 - 12 = 40

For any first card any second card can be matched, so, there are totally 12 * 40 = 480 different combinations